the length of 5000 corn cobs produced from a local wholesaler are measured and recorded. the lengths X, are found to be normally distributed with a mean of 9cm and a variance of 0.25cm². Find the probability that a corn cob selected at random from the wholesaler has a length between 8.5cm and 9.25cm.


1
Expert's answer
2021-10-21T12:03:22-0400

"X\sim N(\mu, \sigma^2)"

Given "\mu=9\ cm , \sigma^2=0.25\ cm^2."


"P(8.5<X<9.25)=P(X<9.25)-P(X\leq8.5)"

"=P(Z<\dfrac{9.25-9}{0.5})-P(Z\leq\dfrac{8.5-9}{0.5})"

"=P(Z<0.5)-P(Z\leq-1)"

"\approx0.691462-0.158655\approx0.532807"

The probability that a corn cob selected at random from the wholesaler has a length between 8.5cm and 9.25cm is "0.5328."